Solve 5 / 6 – 1 / 3 Solve 5 / 6 – 1 / 3 Solve 5 / 6 – 1 / 3

Solve
5 / 6 – 1 / 3
Solve
5 / 6 – 1 / 3
Solve
5 / 6 – 1 / 3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

5 / 6–1 / 3
Taking LCM 6
= [(5 × 1)–(1 × 2)] / 6
= (5–2) / 6
=3 / 6 =1 / 2

Thanks